A car and a motorcycle start from rest at the same time on a straight track, but the motorcycle is 25.0 m behind the car. The car accelerates at a uniform rate of 3.70 m/s2, and the motorcycle at a uniform rate of 4.40 m/s2. How much time elapses before the motorcycle overtakes the car? How far will each have traveled during that time? How far ahead of the car will the motorcycle be 2.00 s later?What is the distance traveled of both car and the motorcycle within that time? (Both vehicles are still accelerating.)
1
Expert's answer
2011-07-29T14:17:41-0400
a1=3.7a2=4.4
h=25
Let t - time elapse before the motorcycle overtakes the car then
h+a1*t2/2=a2*t2/2 = > t=sqrt( 2*h/(a2-a1) )=8.45 sx1=a1*t2/2=132.1 my1=a2*t2/2=157.1 mafter T=t+2=10.45 seconds the car will be far ahead of the motorcycle:

h1=(h+a1*T2/2-a2*T2/2 )=63.23 mx2=a1*T2/2=202.1 my2=a2*T2/2=240.3 m